Grade 12th passPhysical Chemistry

a partical A of certain velocity and certain mass has wavelenth 1Ao.B has mass 25% velocity75% of that A what is its debrogli wavelength

kabeer
10 Years agoGrade 12th pass

1 Answer

Bhavya
10 Years ago
Let velocity of A = v
Let mass of A = m
Wavelength of A =\lambdaA= 1 Angstorm
\lambdaA= h/mv
Mass of B = 25% of m = m/4
Velocity of A = 75% of v = (3/4)*v
\lambdaA/\lambdaB= (h/mv)*(m/4*3v/4)/h
Putting\lambdaA= 1 Angstorm
\lambdaB= (16/3)*1 Angstorm